# Reducing classical communication costs in multiplexed quantum repeaters using hardware-aware quasi-local policies

### v2

@@ Abstract
Future quantum networks will have nodes equipped with multiple quantum memories, allowing for multiplexing and entanglement distillation strategies in order to increase fidelities and reduce waiting times for end-to-end entanglement distribution. In this work, we introduce quasi-local policies for multiplexed quantum repeater chains. In fully-local policies, nodes make decisions based only on knowledge of their own states. In our quasi-local policies, nodes have increased knowledge of the state of the repeater chain, but not necessarily full, global knowledge. Our policies exploit the observation that for most decisions the nodes have to make, they only need to have information about the connected region of the chain they belong to, and not the entire chain. In this way, we not only obtain improved performance over local policies, but we reduce the classical communication (CC) costs inherent to global-knowledge policies. Our policies also outperform the well-known and widely studied nested purification and doubling swapping policy in practically relevant parameter regimes. We also carefully examine the role of entanglement distillation. Via analytical and numerical results, we identify the parameter regimes in which distillation makes sense and is useful. In these regimes, we also address the question: “Should we distill before swapping, or vice versa?” Finally, to provide further practical guidance, we propose an experimental implementation of a multiplexing-based repeater chain, and experimentally demonstrate the key element, a high-dimensional biphoton frequency comb. We then evaluate the anticipated performance of our multiplexing-based policies in such a real-world network through simulation results for two concrete memory platforms, namely rare-earth ions and diamond vacancies.
I Introduction
@@
The amount of classical communication in a network is tied to the amount of knowledge nodes have of each other’s states when making their decisions regarding elementary link generation and entanglement swapping. Indeed, the more knowledge nodes need to make use of, the more classical communication is required to attain that knowledge. The decisions that nodes make, and the amount of knowledge needed to make the decisions, is dictated by the policy being used. In particular, if a policy calls for the nodes to make use of full, global knowledge of the network—meaning that every node should know the state of every other node at all times—then classical communication over the entire span of the network is required in every time step of the protocol. On the other extreme, for fully local policies, in which nodes need to have knowledge only of their own states, only one round of end-to-end classical communication is required at some prescribed final time. In the middle lie quasi-local policies, which we introduce in this work. These policies use to their advantage the realization that nodes only need access to knowledge of the connected portion of the chain they are part of. Disconnected regions of the network can determine their actions independently. Thus, barring only a few time steps when long links are present in the network, spanning lengths on the order of the size of the network, full end-to-end classical communication is superfluous and wasteful. To the best of our knowledge, a thorough analysis of end-to-end waiting times and fidelities, as a function of the amount of knowledge nodes have of each other’s states, has not been conducted. Recent works [44, 45] have highlighted the importance of global knowledge in improving waiting times and fidelities, but without fully accounting for the cost of classical communication. We are therefore motivated to ask our first question:
@@
In this work, we consider this question in the context of entanglement distribution in linear networks (also called “repeater chains”) with multi-memory quantum repeaters and multiplexing [16, 17, 60, 18, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]; see Fig. 1. In the multiplexing approach, simultaneous elementary link generation attempts can be made, and when performing entanglement swapping [5, 78], it is possible to go beyond the trivial parallel approach and perform cross-channel entanglement swapping.
Many of the aforementioned works on multi-memory quantum repeaters with multiplexing have made use of the hardware-agnostic “nested purification and doubling swapping” protocol devised in the seminal works [16, 17] on quantum repeaters. Keeping in mind our goal of developing hardware-aware policiesg, we therefore ask:
@@
In the multiplexing approach, there is also the possibility to perform entanglement purification/distillation [79, 80, 81, 82]. Distillation introduces addition classical communication overhead. We are therefore motivated to ponder whether the potential benefits of distillation on end-to-end fidelities are worth the additional classical communication overhead. Specifically, we ask:
@@
Figure 1: (a) In the multiplexing approach to entanglement distribution, it is possible to perform cross-channel elementary link generation and entanglement swapping. (b) Elementary links are created by multiple entanglement generation sources associated with every channel that connects different pairs of quantum memories. (c) Non-neighbouring nodes are connected by virtual links generated by performing entanglement swapping operations between neighbouring active quantum memories. (d) A linear chain of quantum repeaters with multiple quantum memories capable of establishing multiple physical links, entanglement distillation and entanglement swapping.
I.1 Overview of results
We address the three questions above by extending the framework from Ref. [45] for modeling quantum repeater chains in terms of Markov decision processes to the case of multi-memory quantum repeater chains (see Sec. II). Our main results and contributions comprise four parts.
@@ I.1 Overview of results
    In Sec. III, we define our quasi-local multiplexing-based policies. These policies are generalizations of the so-called “swap-as-soon-as-possible” (swap-asap) policy [30, 31] to the multiplexing setting. We refer to them as strongest neighbor (SN) swap-asap, farthest neighbor (FN) swap-asap, and random swap-asap. These policies are distinguished by how they pair memories within nodes for the purposes of entanglement swapping. Intuitively, the SN swap-asap policy prioritizes the creation of high-fidelity end-to-end links, and FN swap-asap prioritizes lowering the waiting time for the creation of end-to-end links. These policies are also “quasi-local”, requiring some knowledge of the network state but not full, global knowledge.
@@
    In a realistic entanglement distribution setting with imperfect memories, considering classical communication (CC) costs is imperative in order to provide a realistic performance analysis. We incorporate the CC costs associated with our policies, and thereby address (Q1). In particular, we show in Sec. IV that our multiplexing-based FN and SN swap-asap policies outperform fully-local policies that do not require any CC. We thus show that the benefit gained by acquiring information about the network’s state is retained even when the communication costs of acquiring such information is accounted for.
@@
    We also add entanglement distillation to our FN and SN swap-asap policies, and in Sec. V we address (Q3). We find that if the fresh elementary links have high fidelity, using entanglement distillation turns out to be detrimental. On the other hand, when fresh elementary links have low fidelity, and more generally in resource-constrained scenarios, entanglement distillation can be more useful. Nonetheless, even in such cases, we observe examples in which distillation can in fact be detrimental.
    In terms of the relative ordering of distillation and swapping, if links are have low generation probability or low initial fidelity, then distilling first is more advantageous. When the elementary link generation probability is high, it is better to swap first and then distill.
@@
    In Sec. VI, we outline a proof-of-principle experimental realization of a quantum network equipped with multiplexing policies using a high-dimensional bi-photon frequency comb (BFC), and we map the relevant experimental parameters to their counterpart parameters within our modeling framework. Then, in Sec. VI.2, we assess the performance of our multiplexing-based policies within these parameter regimes for two quantum memory platforms, diamond vacancy and rare-earth metal ions. In particular, we consider the number of repeater nodes needed to achieve a desired end-to-end waiting time and fidelity over a distance of 100 km.
II Linear chain quantum network with multiple channels
@@
  • Elementary and virtual links: Entanglement sources establish physical/elementary links between nearest-neighbour quantum memories with probability $p_{\ell}\in[0,1]$. A probabilistic Bell state measurement creates a virtual link between distant quantum memories with success probability $p_{sw}\in[0,1]$. These values are fundamentally limited by the technology used to create the system, e.g., linear optics makes $p_{sw}\leq 0.5$, and $p_{\ell}$ is determined by properties of the source, loss characteristics of the sources, etc. These considerations are relevant for our proposed experimental implementation in Sec. VI, and we provide details on how to determine $p_{\ell}$ and $m^{\star}$ from physical properties of the system in Sec. VI.1.
    We keep track of the fidelity of both elementary and virtual links by assigning every link an age. The age $m$ of a link starts from $m=m_{0}$, when it is first created, with $m_{0}$ being a measure of its initial fidelity with respect to a perfect Bell state. The age then increases in discrete steps, such that $m\in\{0,1,2,3...\}$. Once the age of the link reaches $m^{\star}$, the link is discarded. Throughout this work, we consider a particular Pauli noise model for decoherence of the memories (see Appendix B), such that the fidelity $f(m)$ of a link that has age $m$ is given by
@@
  • Entanglement distillation: Multiple low-fidelity links between any two nodes can be distilled with probability $p_{ds}\in[0,1]$ to form a higher-fidelity link. Throughout this work, we consider the BBPSSW protocol [79], which allows for distillation of two entangled links to one. We provide a summary of the protocol in Appendix C. Briefly, if $f_{1}$ and $f_{2}$ are the fidelities of the two links being distilled, then the success probability $P_{\text{distill}}(f_{1},f_{2})$ of the protocol and the resulting fidelity $F_{\text{distill}}(f_{1},f_{2})$ are given by
@@
    We show in Appendix C that the age after distillation with the BBPSSW protocol is given by the following formula:
@@
    Check the number of active links to the left and right of every node, except for the end nodes. If more than one link on either side is active, rank them based on the policy, as described in detail in Sec. III below. If the doubling policy is being used, pair the links for entanglement swapping only if they are the same length. Perform the entanglement swapping operations in the ranked order. Entanglement swapping is attempted only if the sum of the ages of the two links is less than $m^{\star}$.
@@
    If entanglement distillation is part of the policy, attempt to distill all links between the same two nodes until either one or no link survives; see Sec. III for details. If the policy calls for distilling first, then swapping, interchange steps 2 and 3.
@@
The time steps that we consider can be converted into physical times, and also the ages can be converted to fidelity values, by incorporating the network hardware and design parameters. Examples of such a translation are given in Sec. VI.1 and further details are provided in Appendix B.
II.1 Classical communication
@@
  • A local approach, where all elementary link generation and entanglement swapping attempts are made agnostic to the success or failure of prior swaps. This would then mean that all elementary links (once generated successfully) are retained up to the cutoff time. When an entanglement swap has failed, resulting in both participating links to become inactive, the surrounding nodes must act as if the link is still active until it has reached its expected cutoff. All the classical communication about successful swaps is then done at the end of the protocol. See Appendix F for details about how the end of the protocol is established. Such an approach is fully local and therefore all policy decisions are taken by individual nodes, without any collaboration or exchange of information. The benefit is that no classical communication overheads exist except for heralding of elementary links, and therefore every time step has duration equal to the heralding time only. The cost is instead paid by having to wait for a link to reach its cutoff time, even though it might have become inactive long before that. Distillation can be only included for freshly generated elementary links in a fully local scheme, since as links get older their lengths cannot be predicted locally (since results of swaps are not known).
@@
  • We approximate the CC time by adding $t_{cc}$ time steps to every network evolution step, where $t_{cc}$ is equal to the length of the longest link (number of nodes) involved in any entanglement swap in that step times the CC time between two adjacent nodes of the chain. This allows enough time for the classical communication for all entanglement swaps attempted in that MDP step, and at the end of such a CC-accounted MDP step, all nodes are aware of which nodes they are connected to and the ages of those links (which can also be classically communicated). In most cases this estimate acts as an upper bound to the CC cost, since in practice nodes could perform the next required action as soon as they receive the CC tagged for or relevant to them. See Appendix F for details.
@@
Figure 2: (a) Illustrative example of the FN swap-asap policy. Links on both sides of the $k^{\text{th}}$ node are ranked separately on the basis of their lengths. The longest link is given a rank of 1, while shortest is ranked 3. For entanglement swapping, we thus give first preference to links of rank 1, which would result in the longest possible link. If entanglement swapping fails, then next preference is given to links of rank 2, and so on. (b) In the case of the SN swap-asap policy, link rankings are based on their ages. Preference is given to links with lower ages.
III Quasi-local multiplexing policies
@@
      The Farthest Neighbor (FN) swap-asap policy: This policy prioritizes the creation of virtual links between faraway nodes. Accordingly, we rank the links based on their lengths (farthest links are ranked first), and then entanglement swapping is performed by pairing the links with the same rank. This policy is illustrated with an example in Fig. 2. Intuitively, this policy prioritizes minimizing the average waiting time.
@@
      The Strongest Neighbor (SN) swap-asap policy: This policy prioritizes the creation of virtual links with high fidelity. Accordingly, we rank the links based on their ages (links with the lowest age are ranked first), and then entanglement swapping is performed in that order. Intuitively, this policy aims to create the strongest or highest-fidelity links between the end nodes of the network.
@@
    How should we distill multiple links to one? For this, we introduce the distill-as-soon-as-possible (distill-asap) policy. This is a sequential policy, which can be thought of as a hybrid of the banded and pumping policies [17] (see also Appendix C), and very similar in spirit to the greedy policy defined in Ref. [84]. We sort the links in increasing order of their ages, and then pair them in increasing order. After one round of pairwise distillation, successfully distilled links are paired again and distillation is attempted again, until only one distilled link or no active link remains. This approach therefore uses the benefits of the banded policy, by pairing links of very similar ages, and at the same time, by realising the restrictions introduced by decoherence and finite number of channels, by pumping fresh links into the queue of available links as soon as possible.
@@
The motivation behind our definitions of the FN and SN swap-asap policies is based on prior works [44, 43, 45], which have shown that policies that make use of global knowledge of the state of the repeater chain offer an advantage in terms of waiting time and fidelity. At the same time, such global knowledge comes with considerable classical communication cost and hence is not practical for networks with a large number of nodes and long link distances. This limitation motivated the proposal of so-called quasi-local policies in Ref. [45], which are policies that use global knowledge and coordination between small groups of nodes, but still use local policies such as swap-asap to connect such groups to each other. The SN and FN swap-asap policies being defined here are also quasi-local policies: the nodes do not have knowledge of the entire network, but they have information about other nodes they are connected to. We stress here that access to such information comes with classical communication (CC) costs, but such costs will be much smaller than fully global policies, in which each node knows about the state of every other network node at all times. More precisely, for fully global policies the CC cost in each time step will be proportional to the communication time from one end of the network to another. This would place considerable demands on the coherence times of quantum memories used in the network. On the other extreme, for fully-local policies all CC can be deferred to one round of end-to-end communication at the end of the entanglement distribution protocol. Quasi-local policies try to find a middle ground between these extremes.
@@
We demonstrate via simulations that our quasi-local SN and FN swap-asap multiplexing policies can outperform (in terms of waiting time and end-to-end fidelity) the non-multiplexed, parallel swap-asap policy (as we would expect), and also the random swap-asap policy. We present these results in Appendix A. In the next section, we present results that demonstrate, more importantly, that our policies can outperform fully local multiplexing policies, as well as the well-known and widely-used nested-purification-and-doubling protocol.
@@
IV Performance evaluation
We now evaluate the performance of our policies from Sec. III with respect to the average waiting time and average age of the end-to-end link.
Figure 3: Average waiting time (top) and average age of an end-to-end link (bottom) as a function of the elementary link success probability when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy. The FN swap-asap policy provides smaller values for both figures of merit. Similar trends can also be seen for the SN swap-asap policy.
@@ IV Performance evaluation
Figure 4: (Top) Average waiting time as a function of the number of nodes when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy, and we find that the FN swap-asap policy provides smaller waiting times. More surprisingly, the advantage grows with an increasing number of nodes. This provides evidence of the scalability of the advantage gained by using knowledge of the network state. (Bottom) Number distribution of lengths of entanglement swaps in the FN swap-asap policy. We observe an exponential suppression of entanglement swaps of the order of the length of the chain, which is an important reason for the retained advantage of this policy, even with classical communication included.
@@
A more surprising observation is the fact that even as the number of nodes increases, the advantage of FN swap-asap over the local age-based ranking policy remains strong—in fact, the advantage increases with increasing number of nodes. In Fig. 4 (top), we show the average waiting times for the local age-based ranking and the FN swap-asap policies as a function of the number of nodes. Intuitively, we might have guessed that as the number of nodes increases, so too would the CC cost, because longer and longer swaps would be required to distribute entanglement between the end points of the chain, making quasi-local policies worse off in terms of average waiting times. It is therefore surprising that not only a large improvement using the FN and SN swap-asap policies is obtained for long chains, but further that the improvement increases with the growing scale (number of nodes) of the network.
@@
Nonetheless, we can explain this non-intuitive result as follows. First, we remind the reader that although fully-local policies have the advantage of no CC overheads, they suffer in terms of average waiting time due to the agnostic entanglement swapping. Indeed, waiting to restart links that are already dead is detrimental to the average waiting time. Second, because we use quasi-local policies, the CC costs in a given time step are upper-bounded by the length of the longest entanglement swapping operation, rather than the length of the entire repeater chain. (Recall that we define the length of an entanglement swapping operation as the length, in terms of number of nodes, of the longer of two links involved in the swap.) In a typical time-evolution of a network, from completely disconnected to an end-to-end connected link, for the FN swap-asap policy the number of entanglement swapping operations of length equal to the number of nodes is roughly $O(1)$, i.e., roughly constant with respect to the number of nodes. This is seen in Fig. 4 (bottom). The average number distribution of the length of entanglement swaps shows an exponentially suppressed tail. Therefore, there are very few time steps in which the CC cost is given by the entire length of the repeater chain. This justifies and supports the use of quasi-local policies over global policies for large networks, because in the latter, the CC cost is equal for each time step and given by the entire length of the repeater chain, while for quasi-local policies the CC cost can be considerably less. Also see Appendix F for details.
IV.1 Comparison with doubling
@@
Figure 5: Comparison of (top) average waiting times and (bottom) average age of the end-to-end link for the FN and SN swap-asap policies and the SN doubling policy. FN swap-asap outperforms both SN swap-asap and SN doubling in terms of average waiting time, and SN swap-asap performs the best in terms of average age of the end-to-end link, for all elementary link success probabilities ($p_{\ell}$).

### v3

@@ Abstract
Future quantum networks will have nodes equipped with multiple quantum memories, allowing for multiplexing and entanglement distillation strategies for long-distance entanglement distribution. In this work, we focus on quasi-local policies for multiplexed quantum repeater chains. In fully-local policies, nodes use the knowledge of only their own states, whereas more efficient global policies use knowledge of the entire network state. The classical communication costs of using this knowledge have not been explored in existing literature. We show that quasi-local policies not only obtain improved performance over local policies, but also reduce classical communication costs considerably. Our policies also outperform the widely studied nested purification and doubling policy in practical parameter regimes. We identify parameter regimes where distillation is useful and address the question: “Should we distill before swapping, or vice versa?” Finally, we propose an implementation scheme for a multiplexed repeater chain, experimentally demonstrate the key element, a high-dimensional biphoton frequency comb, and evaluate its anticipated performance using our multiplexing-based policies.
Introduction
@@
The amount of classical communication in a network is tied to the amount of knowledge nodes have of each other’s states when making their decisions regarding elementary link generation and entanglement swapping. Indeed, the more knowledge nodes need to make use of, the more classical communication is required to attain that knowledge. The decisions that nodes make, and the amount of knowledge needed to make the decisions, is dictated by the policy being used. In particular, if a policy calls for the nodes to make use of full, global knowledge of the network—meaning that every node should know the state of every other node at all times—then classical communication over the entire span of the network is required in every time step of the protocol. On the other extreme, for fully local policies, in which nodes need to have knowledge only of their own states, only one round of end-to-end classical communication is required at some prescribed final time. In the middle lie quasi-local policies, which we consider in this work. These policies use to their advantage the realization that nodes only need access to knowledge of the connected portion of the chain they are part of. Disconnected regions of the network can determine their actions independently. Thus, barring only a few time steps when long links are present in the network, spanning lengths on the order of the size of the network, full end-to-end classical communication is superfluous and wasteful. Previous works [60, 61, 62] have considered quasi-local policies in the context of continuous entanglement distribution, while we consider on-demand entanglement distribution. To the best of our knowledge, a thorough analysis of end-to-end waiting times and fidelities, as a function of the amount of knowledge nodes have of each other’s states, and taking classical communication costs into account, has not been conducted. Recent works [44, 45] have highlighted the importance of global knowledge in improving waiting times and fidelities, but without fully accounting for the cost of classical communication. We are therefore motivated to ask our first question:
@@
In this work, we consider this question in the context of entanglement distribution in linear networks (also called “repeater chains”) with multi-memory quantum repeaters and multiplexing [16, 17, 63, 18, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 34, 76, 77, 78, 79, 80, 81]; see Fig. 1. In the multiplexing approach, simultaneous elementary link generation attempts can be made, and when performing entanglement swapping [5, 82], it is possible to go beyond the trivial parallel approach and perform cross-channel entanglement swapping.
Many of the aforementioned works on multi-memory quantum repeaters with multiplexing have made use of the hardware-agnostic “nested purification and doubling swapping” protocol devised in the seminal works [16, 17] on quantum repeaters. Keeping in mind our goal of developing hardware-aware policies, we therefore ask:
@@
In the multiplexing approach, there is also the possibility to perform entanglement purification/distillation [83, 84, 85, 86]. Distillation introduces additional classical communication overhead. We are therefore motivated to ponder whether the potential benefits of distillation on end-to-end fidelities are worth the additional classical communication overhead. Specifically, we ask:
@@
Figure 1: Multiplexed entanglement distribution. (a) In the multiplexing approach to entanglement distribution, it is possible to perform cross-channel elementary link generation and entanglement swapping. (b) Elementary links are created by multiple entanglement generation sources associated with every channel that connects different pairs of quantum memories. (c) Non-neighboring nodes are connected by virtual links generated by performing entanglement swapping operations between neighboring active quantum memories. (d) A linear chain of quantum repeaters with multiple quantum memories capable of establishing multiple physical links, entanglement distillation and entanglement swapping.
In this paper, we address the three questions above by extending the framework from Ref. [45] for modeling quantum repeater chains in terms of Markov decision processes (MDPs) to the case of multi-memory quantum repeater chains. We define quasi-local multiplexing-based policies, which are generalizations of the so-called “swap-as-soon-as-possible” (swap-asap) policy [30, 31], to the multiplexing setting. For near-term quantum networks with imperfections such as memories with limited coherence time, probabilistic entanglement swapping, etc., classical communication (CC) costs are considerable, but usually ignored in existing literature. We incorporate the CC costs associated with our policies, and show that these quasi-local policies outperform fully-local policies that do not require any CC. We also show that our policies can outperform the doubling policy in practically relevant parameter regimes. Next, we also add entanglement distillation to our model, and address several policy questions. Finally, we consider a proof-of-principle experimental realization of a multiplexed quantum network equipped using a high-dimensional bi-photon frequency comb (BFC) and assess its performance.
Results and discussion
Overview of results
Our main results and contributions comprise four parts.
@@
    We define our quasi-local multiplexing-based policies. These policies are generalizations of the so-called “swap-as-soon-as-possible” (swap-asap) policy [30, 31] to the multiplexing setting. We refer to them as strongest neighbor (SN) swap-asap, farthest neighbor (FN) swap-asap, and random swap-asap. These policies are distinguished by how they pair memories within nodes for the purposes of entanglement swapping. Intuitively, the SN swap-asap policy prioritizes the creation of high-fidelity end-to-end links, and FN swap-asap prioritizes lowering the waiting time for the creation of end-to-end links. These policies are also “quasi-local”, requiring some knowledge of the network state but not full, global knowledge.
@@
    In a realistic entanglement distribution setting with imperfect memories, considering classical communication (CC) costs is imperative in order to provide a realistic performance analysis. In Performance evaluation, we incorporate the CC costs associated with our policies, and thereby address (Q1). In particular, we show that our multiplexing-based FN and SN swap-asap policies outperform fully-local policies that do not require any CC. We thus show that the benefit gained by acquiring information about the network’s state is retained even when the communication costs of acquiring such information is accounted for.
@@
    In Distillation-based policies, we also add entanglement distillation to our FN and SN swap-asap policies, and then address (Q3). We find that if the fresh elementary links have high fidelity, using entanglement distillation turns out to be detrimental. On the other hand, when fresh elementary links have low fidelity, and more generally in resource-constrained scenarios, entanglement distillation can be more useful. Nonetheless, even in such cases, we observe examples in which distillation can in fact be detrimental.
    In terms of the relative ordering of distillation and swapping, if links have low generation probability or low initial fidelity, then distilling first is more advantageous. When the elementary link generation probability is high, it is better to swap first and then distill.
@@
    In Proposed Experimental implementation, we outline a proof-of-principle experimental realization of a quantum network equipped with multiplexing policies using a high-dimensional bi-photon frequency comb (BFC), and we map the relevant experimental parameters to their counterpart parameters within our modeling framework. Then we assess the performance of our multiplexing-based policies within these parameter regimes for two quantum memory platforms, diamond vacancy and rare-earth metal ions. In particular, we consider the number of repeater nodes needed to achieve a desired end-to-end waiting time and fidelity over a distance of 100 km.
Linear chain quantum network with multiple channels
@@
  • Elementary and virtual links: Entanglement sources establish physical/elementary links between nearest-neighbor quantum memories with probability $p_{\ell}\in[0,1]$. A probabilistic Bell state measurement creates a virtual link between distant quantum memories with success probability $p_{sw}\in[0,1]$. These values are fundamentally limited by the technology used to create the system, e.g., linear optics makes $p_{sw}\leq 0.5$, and $p_{\ell}$ is determined by properties of the source, loss characteristics of the sources, etc. These considerations are relevant for our proposed experimental implementation, and we also provide details on how to determine $p_{\ell}$ and $m^{\star}$ from physical properties of the system.
    We keep track of the fidelity of both elementary and virtual links by assigning every link an age. The age $m$ of a link starts from $m=m_{0}$, when it is first created, with $m_{0}$ being a measure of its initial fidelity with respect to a perfect Bell state. The age then increases in discrete steps, such that $m\in\{0,1,2,3...\}$. Once the age of the link reaches $m^{\star}$, the link is discarded. Throughout this work, we consider a particular Pauli noise model for decoherence of the memories (see “Methods”), such that the fidelity $f(m)$ of a link that has age $m$ is given by
@@
  • Entanglement distillation: Multiple low-fidelity links between any two nodes can be distilled with probability $p_{ds}\in[0,1]$ to form a higher-fidelity link. Throughout this work, we consider the BBPSSW protocol [83], which allows for distillation of two entangled links to one. We provide a summary of the protocol in “Methods”. Briefly, if $f_{1}$ and $f_{2}$ are the fidelities of the two links being distilled, then the success probability $P_{\text{distill}}(f_{1},f_{2})$ of the protocol and the resulting fidelity $F_{\text{distill}}(f_{1},f_{2})$ are given by
@@
    We show in “Methods” that the age after distillation with the BBPSSW protocol is given by the following formula:
@@
    Check the number of active links to the left and right of every node, except for the end nodes. If more than one link on either side is active, rank them based on the policy, as described in detail in Quasi-local multiplexing policies below. If the doubling policy is being used, pair the links for entanglement swapping only if they are the same length. Perform the entanglement swapping operations in the ranked order. Entanglement swapping is attempted only if the sum of the ages of the two links is less than $m^{\star}$.
@@
    If entanglement distillation is part of the policy, attempt to distill all links between the same two nodes until either one or no link survives (also see Supplementary Note 2). If the policy calls for distilling first, then swapping, interchange steps 2 and 3.
@@
The time steps that we consider can be converted into physical times, and also the ages can be converted to fidelity values, by incorporating the network hardware and design parameters. Examples of such a translation are given in Proposed experimental implementation and further details are provided in “Methods”.
Classical communication
@@
  • A local approach, where all elementary link generation and entanglement swapping attempts are made agnostic to the success or failure of prior swaps. This would then mean that all elementary links (once generated successfully) are retained up to the cutoff time. When an entanglement swap has failed, resulting in both participating links to become inactive, the surrounding nodes must act as if the link is still active until it has reached its expected cutoff. All classical communication about successful swaps is then done at the end of the protocol. See “Methods” for details about how the end of the protocol is established. Such an approach is fully local and therefore all policy decisions are taken by individual nodes, without any collaboration or exchange of information. The benefit is that no classical communication overheads exist except for heralding of elementary links, and therefore every time step has duration equal to the heralding time only. The cost is instead paid by having to wait for a link to reach its cutoff time, even though it might have become inactive long before that. Distillation can be only included for freshly generated elementary links in a fully local scheme, since as links get older their lengths cannot be predicted locally (since results of swaps are not known).
@@
  • We approximate the CC time by adding $t_{cc}$ time steps to every network evolution step, where $t_{cc}$ is equal to the length of the longest link (number of nodes) involved in any entanglement swap in that step times the CC time between two adjacent nodes of the chain. This allows enough time for the classical communication for all entanglement swaps attempted in that MDP step, and at the end of such a CC-accounted MDP step, all nodes are aware of which nodes they are connected to and the ages of those links (which can also be classically communicated). In most cases this estimate acts as an upper bound to the CC cost, since in practice nodes could perform the next required action as soon as they receive the CC tagged for or relevant to them. See “Methods” for details.
@@
Figure 2: Our proposed entanglement swapping policies for multiplexed repeater chains. (a) Illustrative example of the FN swap-asap policy. Links on both sides of the $k^{\text{th}}$ node are ranked separately on the basis of their lengths. The longest link is given a rank of 1, while shortest is ranked 3. For entanglement swapping, we pair links of rank 1, which would result in the longest possible link. Next links of rank 2 are connected, and so on. (b) In the case of the SN swap-asap policy, link rankings are based on their ages.
Quasi-local multiplexing policies
@@
      The Farthest Neighbor (FN) swap-asap policy: This policy prioritizes the creation of virtual links between faraway nodes. Accordingly, we rank the links based on their lengths (farthest links are ranked first), and then entanglement swapping is performed by pairing the links with the same rank. This policy is illustrated with an example in Fig. 2(a). Intuitively, this policy prioritizes minimizing the average waiting time.
@@
      The Strongest Neighbor (SN) swap-asap policy: This policy prioritizes the creation of virtual links with high fidelity. Accordingly, we rank the links based on their ages (links with the lowest age are ranked first), and then entanglement swapping is performed by pairing links with same rank; see Fig. 2(b). Intuitively, this policy aims to create the strongest or highest-fidelity links between the end nodes of the network.
@@
    How should we distill multiple links to one? For this, we introduce the distill-as-soon-as-possible (distill-asap) policy. This is a sequential policy, which can be thought of as a hybrid of the banded and pumping policies [17] (see Supplementary Note 2 for details), and very similar in spirit to the greedy policy defined in Ref. [88]. We sort the links in increasing order of their ages, and then pair them in increasing order. After one round of pairwise distillation, successfully distilled links are paired again and distillation is attempted again, until only one distilled link or no active link remains. This approach therefore uses the benefits of the banded policy, by pairing links of very similar ages, and at the same time, by realizing the restrictions introduced by decoherence and finite number of channels, by pumping fresh links into the queue of available links as soon as possible.
@@
We demonstrate via Monte Carlo simulations of the underlying Markov decision process that our quasi-local SN and FN swap-asap multiplexing policies can outperform (in terms of waiting time and end-to-end fidelity) the non-multiplexed, parallel swap-asap policy (as we would expect), and also the random swap-asap policy. We present these results in Supplementary Note 1 (also see Supplementary Note 4 for some alternative quasi-local multiplexing policies). In the next section, we present results that demonstrate, more importantly, that our policies can outperform fully local multiplexing policies, as well as the well-known and widely-used nested-purification-and-doubling protocol.
@@
Performance evaluation
We now evaluate the performance of our policies described above with respect to the average waiting time and average age of the end-to-end link. All of our results were obtained using Monte Carlo simulations, and details about the simulations can be found in “Methods”.
Figure 3: Quasi-local policies outperform fully-local policies. Average waiting time (a) and average age of an end-to-end link (b) as a function of the elementary link success probability when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy. The FN swap-asap policy provides smaller values for both figures of merit. Similar trends can also be seen for the SN swap-asap policy.
@@
Figure 4: Advantage of quasi-local policies is scalable and grows with increasing number of nodes. (a) Average waiting time as a function of the number of nodes when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy, and we find that the FN swap-asap policy provides smaller waiting times. More surprisingly, the advantage grows with an increasing number of nodes. This provides evidence of the scalability of the advantage gained by using knowledge of the network state. (b) Number distribution of lengths of entanglement swaps in the FN swap-asap policy. We observe an exponential suppression of entanglement swaps of the order of the length of the chain, which is an important reason for the retained advantage of this policy, even with classical communication included.
@@
A more surprising observation is the fact that even as the number of nodes increases, the advantage of FN swap-asap over the local age-based ranking policy remains strong—in fact, the advantage increases with increasing number of nodes. In Fig. 4(a), we show the average waiting times for the local age-based ranking and the FN swap-asap policies as a function of the number of nodes. Intuitively, we might have guessed that as the number of nodes increases, so too would the CC cost, because longer and longer swaps would be required to distribute entanglement between the end points of the chain, making quasi-local policies worse off in terms of average waiting times. It is therefore surprising that not only a large improvement using the FN and SN swap-asap policies is obtained for long chains, but further that the improvement increases with the growing scale (number of nodes) of the network.
@@
Nonetheless, we can explain this non-intuitive result as follows. First, we remind the reader that although fully-local policies have the advantage of no CC overheads, they suffer in terms of average waiting time due to the agnostic entanglement swapping. Indeed, waiting to restart links that are already dead is detrimental to the average waiting time. Second, because we use quasi-local policies, the CC costs in a given time step are upper-bounded by the length of the longest entanglement swapping operation, rather than the length of the entire repeater chain. (Recall that we define the length of an entanglement swapping operation as the length, in terms of number of nodes, of the longer of two links involved in the swap.) In a typical time-evolution of a network, from completely disconnected to an end-to-end connected link, for the FN swap-asap policy the number of entanglement swapping operations of length equal to the number of nodes is roughly $O(1)$, i.e., roughly constant with respect to the number of nodes. This is seen in Fig. 4(b). The average number distribution of the length of entanglement swaps shows an exponentially suppressed tail. Therefore, there are very few time steps in which the CC cost is given by the entire length of the repeater chain. This justifies and supports the use of quasi-local policies over global policies for large networks, because in the latter, the CC cost is equal for each time step and given by the entire length of the repeater chain, while for quasi-local policies the CC cost can be considerably less.
Comparison with doubling
@@
Figure 5: Strongest Neighbor and Farthest Neighbor swap-asap policies outperform the doubling policy. Comparison of (a) average waiting times and (b) average age of the end-to-end link for the FN and SN swap-asap policies and the SN doubling policy. FN swap-asap outperforms both SN swap-asap and SN doubling in terms of average waiting time, and SN swap-asap performs the best in terms of average age of the end-to-end link, for all elementary link success probabilities ($p_{\ell}$).